Every 3 out of 1000 usb sticks is bad. If an employee randomly selects 10. What is the probability that one or more is bad?
Answer #1

Probability that usb stick is bad, p = 3/1000 = 0.003

Probability that usb stick is not bad, p = 1- 0.003 = 0.997

Sample size, n = 10

Probability that one or more is bad, P(x >=1) = 1- P(none is bad)

1 - (0.997)10

1-0.97040.0296